OpenStudy (kainui):

I need to know the name of a function that's 1 at every integer and 0 everywhere else. Anyone have any info on a function like this?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

they call this kind of thing a "characteristic function"

OpenStudy (misty1212):

this would be the characteristic function of the integers

OpenStudy (kainui):

Should be nice to see, I'll look into that. Here's some alternative definition I came up with among others to help me out in the mean time: \[\Large f(x) = \lim_{r \to \infty} \cos^r(\pi x)\]

OpenStudy (kainui):

@misty1212 Hey could you give me some more information, I'm sorta having trouble in google searching "Characteristc Function" haha

ganeshie8 (ganeshie8):

should be possible with floor/ceil : \[f(x) = -\lfloor x\rfloor -\lfloor -x\rfloor \]

OpenStudy (kainui):

That's something I hadn't thought of that looks pretty cool! The reason I'm interested in this function is because: \[\Large \tau (x) = \sum_{n=1}^{ \infty } f(\frac{x}{n})\]

ganeshie8 (ganeshie8):

that gives 0 when x is integer and 1 otherwise

ganeshie8 (ganeshie8):

(refering to floor function)

OpenStudy (kainui):

Oh so I need 1-that I think? I'm going to play with it and just figure it out one sec.

OpenStudy (kainui):

Here's what I believe/am saying right now. \[\Large \lim_{r \to \infty} \cos^r( \pi x) = 1+ \lfloor x \rfloor + \lfloor -x \rfloor\]

OpenStudy (zarkon):

I typically call them indicator functions http://en.wikipedia.org/wiki/Indicator_function when I hear characteristic function I think of this http://en.wikipedia.org/wiki/Characteristic_function_%28probability_theory%29

OpenStudy (kainui):

Ahhh good and it's related to the Heaviside and Delta functions, thanks this is something I think I can work with now.

OpenStudy (zarkon):

\[\mathbb{1}_{\mathbb{Z}}(x)\]

OpenStudy (kainui):

Interesting, I think this might be the first time I've seen a function denoted with a number instead of a letter before haha.

OpenStudy (misty1212):

what is wrong with plain old \[f(x) = \left\{\begin{array}{rcc} 1 & \text{if} ~x\in\mathbb{Z} \\ 0& \text{otherwise} \end{array} \right. \]

OpenStudy (zarkon):

that would require one to not be lazy

OpenStudy (zarkon):

I'm too lazy to write that much

OpenStudy (misty1212):

how about "one if x is an integer, zero otherwise"?

OpenStudy (zarkon):

That seems like a lot of words to me. I don't like to write that much when I'm on vacation.